If the coefficient of friction is 0.6,calculate the angle of friction 
 

Grade:11

1 Answers

Arun
4 years ago
 
Formula for angle of friction is
\tan \theta = \mu _{s}
Where
Theta = Angle of friction
\mu _{s} = Coefficient of Static Friction
Here taking  \mu _{s}= 0.6
\theta = \arctan 0.6
\theta \approx 30.96